OpenStudy (anonymous):

what are the first 5 terms of the sequence given by the formula An=4n+1? a.1,5,9,13,17 b.9,13,17,21,25 c.5,9,13,17,21 d.4,8,12,16,20

OpenStudy (anonymous):

c

OpenStudy (anonymous):

4(1)+1=5 4(2)+1=9 4(3)+1=13 4(4)+1=17 so its c